# 40 please Question 39 Provide an appropriate response. Data sets A and B are dependent. Find the test statistic. А 6.1 7.1 9.0 6.0 6.1
Answer #1
A B d = A-B (d-d̅)2
13 11 2 49
11 7 4 25
30 8 22 169
26 18 8 1
14 5 9 0
Σd = 45 Σ(d-d̅)2 = 244

d̅ = d/n = 45/5 = 9

s = (Σ(d-d̅)2/(n-1))^0.5 = (244/4)^0.5 = 7.81

Test statistic = d̅ /(s/n^0.5) = 9/(7.81/5)^0.5 = 7.201